OpenStudy (anonymous):

5x-4y=10. write in slope intercept form. is this correct: y=5/4x-15/4 ?

OpenStudy (anonymous):

Where did the fifteen in the numerator come from? I get\[5x-4y=10\implies -4y=-5x+10\implies y=\frac{5}{4}x-\frac{5}{2}\]The constant term simplifies from 10/4.

hero (hero):

\[5x - 4y = 10\]\[-4y = -5x + 10\]\[y = \frac{-5x + 10}{-4}\]\[y = \frac{-5x}{-4} + \frac{10}{-4}\]\[y = \frac{5x}{4} + \frac{5}{2}\]

hero (hero):

-5/2 , sorry

OpenStudy (anonymous):

thx for showing the steps:)